# Quadratic-form Optimal Transport

### v2

@@ Appendix F Omitted proofs of results from Section 4
Therefore, the objective $\mathbb{E}_{\pi\otimes\pi}[c(X,Y,X^{\prime},Y^{\prime})]$ is a linear or quadratic function of $\mathbb{E}_{\pi}[XY]$. On the other hand, the upper and lower bounds for $\mathbb{E}_{\pi}[XY]$ are attained explicitly by $\pi_{\mathrm{com}}$ and $\pi_{\mathrm{ant}}$, respectively. In addition, for any $\beta$ in the interval
$\Big{[}\inf_{\pi\in\Pi(\mu,\nu)}\mathbb{E}_{\pi}[XY],\sup_{\pi\in\Pi(\mu,\nu)}\mathbb{E}_{\pi}[XY]\Big{]},$ (F.2)
there exists $\lambda\in[0,1]$ such that $\mathbb{E}_{\pi_{\rm x}^{\lambda}}[XY]=\beta$ because $\lambda\mapsto\mathbb{E}_{\pi_{\rm x}^{\lambda}}[XY]$ is affine. This shows that a QOT minimizer $\pi_{\lambda}$ exists. The last statement follows by noting that the range (F.2) for $\beta$ is not a singleton, and any $\beta$ in (F.2) can be a unique minimizer for some choices of $C_{1},C_{2},\alpha_{1},\alpha_{2}$ in (F.1), which are arbitrary. Hence, any $\lambda\in[0,1]$ can yield a unique minimizer in the class $(\pi_{\rm x}^{\lambda})_{\lambda\in[0,1]}$. ∎
Proof of Proposition 4.3.
This follows from the same arguments in the proof of Proposition 4.1 by noting that every point in the interval
$\left[\inf_{\pi\in\Pi(\mu,\nu)}\int f\,\mathrm{d}\pi,\sup_{\pi\in\Pi(\mu,\nu)}\int f\,\mathrm{d}\pi\right]$
is attained by some $\pi_{\lambda}$. ∎
@@
Hence, either $X+Y=X’+Y’$ ’ or $X-Y=X’-Y’$ ’ almost surely. Since the two sides are independent, we have either $X+Y$ is a constant (only if $(X,Y)\buildrel\mathrm{law}\over{\sim}\pi_{\mathrm{ant}}$) or $X-Y$ is a constant (only if $(X,Y)\buildrel\mathrm{law}\over{\sim}\pi_{\mathrm{com}}$). Since $\mu=\nu$, the comonotone coupling verifies $X-Y=0$; the antimonotone coupling verifies $X+Y$ is a constant if and only if $\mu$ is symmetric. This completes the proof. ∎
@@
$X\mbox{ is uniformly distributed on }p\mbox{ points in }[n];$
$Y\mbox{ is uniformly distributed on }q\mbox{ points in }[n].$
@@
Let $(X,Y)$ be given by the claimed optimal solution. Suppose on the contrary that there exist independent random variables $\hat{X},\hat{Y}$ satisfying the constraints, whose joint law $(\hat{X},\hat{Y})$ is different from $(X,Y)$, and furthermore,
@@
for some $\gamma>0$. Note that $(\hat{X},\hat{Y})$ is absolutely continuous with bounded density by the constraints. Then there exist a sequence of random variables $(\hat{X}_{n},\hat{Y}_{n})_{n\geqslant 1}$ such that:
@@
subject to $\xi_{1},\xi_{2}\mbox{ are independent r.v.s on }[0,1]\mbox{ with respective densities }f_{\xi_{1}},f_{\xi_{2}};$
@@ Appendix G Omitted proofs of results from Section 5
By (5.5), we have for all $u\in\mathcal{D}$,

### v4

@@ Appendix F Omitted proofs of results from Section 5
We prove Proposition 5.3 first. The proof of Proposition 5.1 uses Proposition 5.3.
Proof of Proposition 5.3.
Using the independence of $(X,Y)$ and $(X^{\prime},Y^{\prime})$, we see that the objective $\mathbb{E}_{\pi\otimes\pi}[c(X,Y,X^{\prime},Y^{\prime})]$ is a linear or quadratic function of $\mathbb{E}_{\pi}[f(X,Y)]$. On the other hand, since $f$ is submodular, the upper and lower bounds for $\mathbb{E}_{\pi}[f(X,Y)]$ are attained explicitly by $\pi_{\mathrm{com}}$ and $\pi_{\mathrm{ant}}$, respectively. In addition, for any $\beta$ in the interval
$\Big{[}\inf_{\pi\in\Pi(\mu,\nu)}\mathbb{E}_{\pi}[f(X,Y)],\sup_{\pi\in\Pi(\mu,\nu)}\mathbb{E}_{\pi}[f(X,Y)]\Big{]},$ (F.1)
there exists $\lambda\in[0,1]$ such that $\mathbb{E}_{\pi_{\rm x}^{\lambda}}[f(X,Y)]=\beta$ because $\lambda\mapsto\mathbb{E}_{\pi_{\rm x}^{\lambda}}[f(X,Y)]$ is affine. This shows that a QOT minimizer $\pi_{\lambda}$ exists. The last statement follows by noting that, for some choices of $C_{1},C_{2},\alpha_{1},\alpha_{2}\in\mathbb{R}$ (which are arbitrary) in (5.2), the range (F.1) for $\beta$ is not a singleton and any $\beta$ in (F.1) can be a unique minimizer. Hence, any $\lambda\in[0,1]$ can yield a unique minimizer in the class $(\pi_{\rm x}^{\lambda})_{\lambda\in[0,1]}$. ∎
@@
The desired result is then a special case of Proposition 5.3 since $(x,y)\mapsto xy$ is supermodular. The last statement follows since any combination of $(C_{1},C_{2},\alpha_{1},\alpha_{2})$ can be achieved by certain choices of the quadratic functions $f,g$. ∎
@@
Hence, either X+Y=X+YX+Y=X’+Y’ or XY=XYX-Y=X’-Y’ almost surely. Since the two sides are independent, we have either $X+Y$ is a constant (only if $(X,Y)\buildrel\mathrm{law}\over{\sim}\pi_{\mathrm{ant}}$) or XYX-Y is a constant (only if $(X,Y)\buildrel\mathrm{law}\over{\sim}\pi_{\mathrm{com}}$). Since $\mu=\nu$, the comonotone coupling verifies XY=0X-Y=0; the antimonotone coupling verifies $X+Y$ is a constant if and only if $\mu$ is symmetric. This completes the proof. ∎
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ Appendix G Omitted proofs of results from Section 6
By (6.5), we have for all $u$ such that $u^{2}\in\mathcal{D}$,